# Learning on LLM Output Signatures for Gray-Box Behavior Analysis

### v2

@@ 5.4 Run-Time Analysis
We conclude this section by discussing our comprehensive training and inference timings, reported in Section B.9 and Table 6. We remark how LOS-Net features an extremely contained detection latency: $\approx 10^{-5}$s per inference fwd-pass. Training is also efficient, typically completing in under one hour on a single NVIDIA L-40 GPU, and often taking significantly shorter. To contextualize this computational efficiency w.r.t. methods relying on multiple prompting/generations [35, 23], we measured the generation time of Mistral-7B-Instruct. Averaging over the first 10 samples from the HotpotQA dataset, we observed an average generation time of $1.93\pm 0.18$ seconds per response on a server equipped with eight NVIDIA L-40 GPUs. The methods in [35] and [23] require 5 and 10 generations per detection, resp., rendering their computation overhead around six orders of magnitude higher than that of LOS-Net.
@@ 6 Conclusions
We proposed LOS-Net, an efficient method to detect data contamination and hallucinations in LLMs by leveraging their output signatures (LOS), defined as the union of Token Distribution Sequences (TDS) and Actual Token Probabilities (ATP). LOS-Net consists of a lightweight transformer with learnable rank encodings applied on the whole LOS. We proved it unifies and extends existing gray-box methods under a general framework, and experimentally showed it outperforms state-of-the-art gray-box methods across datasets and LLMs. It also exhibited strong generalization capabilities of LOS-Net, both across datasets and across LLMs. Our framework could be applied to other tasks, such as detecting LLM-generated content. Additional sources of information can also be incorporated, e.g., in the absence of latency constraints, it can be interesting to include “exact-token” flags as proposed by [35]. Last, the LOS can be extended to account for multiple prompting [23].
Limitations. By operating in a gray-box setting, our approach is widely applicable, but misses access to predictive information residing, e.g., in the LLM’s hidden states. Additionally, we note that sorting the TDS tensor removes word alignment across the vocabulary, which may be limiting in some cases.
Acknowledgments
The authors are grateful to Beatrice Bevilacqua for insightful discussions. G.B. is supported by the Jacobs Qualcomm PhD Fellowship. F.F. conducted this work supported by an Aly Kaufman and an Andrew and Erna Finci Viterbi Post-Doctoral Fellowship. Y.G. is supported by the UKRI Engineering and Physical Sciences Research Council (EPSRC) CDT in Autonomous and Intelligent Machines and Systems (grant reference EP/S024050/1). H.M. is a Robert J. Shillman Fellow and is supported by the Israel Science Foundation through a personal grant (ISF 264/23) and an equipment grant (ISF 532/23). D.L. is funded by an NSF Graduate Fellowship. Research was also supported by the Israeli Ministry of Science, Israel-Singapore binational grant 207606.
@@ Appendix A Proofs
We define $\mathcal{D}\coloneqq\mathcal{X}\times\mathcal{M}.$ Recall that the target function we want to approximate is the gated scoring function $R$ as defined in Equation 5, which can be written as follows:
@@
Define $f^{(1)}:\mathcal{D}\rightarrow\mathbb{R}^{N_{\text{max}}}$ to be the components of the sum in Equation 6:
@@
As a result, the input to the transformer encoder in our architecture (see Equation 3) becomes $\mathbf{X}^{\prime}||\mathbf{p}\in\mathbb{R}^{N_{\text{max}}\times(V_{\text{max}}+1)}$.
This simplifies our architecture in Equation 3 to:
@@
[…]
@@
Next, observe that $f^{(1)}\in L^{1}(\mathcal{D})$. To see this, first note that $f^{(1)}$ is measurable. The indicator function is measurable because the indicator set is the preimage of the measurable function $\kappa(x)-T(x)$ on the closed set $[0,\infty)$. Thus, $f^{(1)}$, being a product of measurable functions, is measurable. Next, we show that the $L^{1}$ norm is finite. This is true because $f^{(1)}$ is a product of the integrable function $g$ and the bounded function $\mathbf{1}$ on the compact domain $\mathcal{D}$.
@@
$\boldsymbol{\sigma}_{i}$ $=\sqrt{\mathbb{E}_{\mathbf{X}^{\prime}_{i}}[(\log(\mathbf{X}^{\prime}_{i})-\boldsymbol{\mu}_{i})^{2}]}=\sqrt{\sum_{v=1}^{V}\mathbf{X}^{\prime}_{i,v}\cdot\big{(}\log(\mathbf{X}^{\prime}_{i,v})-\boldsymbol{\mu}_{i}\big{)}^{2}},$ (22)
Where $\mathbf{X}^{\prime}$ is given from Equation 1.
@@
To prove Corollary 4.3, it suffices to show the following. First (i), that the baselines can be implemented as in Equation 5, given their sequence length and vocabulary size satisfy, $N\leq N_{\text{max}},\quad V\leq V_{\text{max}}$, where values in the inputs for indices larger than $N,V$ are ‘padded’ with e.g., $-1$. Second (ii), that their implementations are realized with $\kappa$, $T$, and $g$ which are all measurable, and with $g$ also integrable.
(i) Let us slightly modify the implementations provided in the Proof for Proposition 4.1 to correctly account for padding values. Let us conveniently define:
$\alpha:\mathbb{R}\rightarrow\mathbb{R},$ $\quad\alpha(x)=1-\mathrm{ReLU}(-x)=\begin{cases}1&x\geq 0\\ 1+x&x<0\end{cases}$
@@
$\kappa(\mathbf{X}^{\prime},\mathbf{p})=-\beta(\mathbf{p})\quad T(\mathbf{X}^{\prime},\mathbf{p})=-\min(\beta(\mathbf{p}))\quad g(\mathbf{X}^{\prime},\mathbf{p})=\mathbf{p}\quad M=2,f\equiv\text{id.}$.
@@
$\kappa(\mathbf{X}^{\prime},\mathbf{p})=-\beta(\mathbf{p})\quad T(\mathbf{X}^{\prime},\mathbf{p})=-\big{(}\text{sort}(\beta(\mathbf{p}))_{\left\lceil\frac{K}{100}\cdot N_{\text{eff}}\right\rceil}\big{)}\quad g(\mathbf{X}^{\prime},\mathbf{p})=\frac{\log(\beta(\mathbf{p}))}{\left\lceil\frac{K}{100}\cdot N_{\text{eff}}\right\rceil}\quad M=2,f\equiv\text{id.}$.
@@
MinK%++. Before illustrating how this baseline is implemented, we note the following. In order for the normalization of log-probs to be well-defined, it is required that: (1) $\mathbf{\mu}$ is finite, (2) the denominator is greater than 00. As for (1), we note that null probability values ($X_{i,v}=0$) would be problematic, as they would cause the $\log$ function to output $\infty$- ∞. We assume, in this case, that all probability values lie in $[\epsilon_{1},1]$, with $\epsilon_{1}$ being a small value such that $10<\epsilon_{1}<$ 1. Regarding (2), we see that the problematic situation would occur in cases where the probability distribution is uniform. We assume to handle this case by adding a small positive constant $\epsilon_{2}>0$ in the denominator, so that the normalization would take the form: $\bar{\mathbf{p}}=\frac{\log(\mathbf{p})-\boldsymbol{\mu}}{\boldsymbol{\sigma}+\epsilon_{2}}$.
@@
$\beta_{1}=\beta(\cdot;2,\text{id.})\quad\beta^{i}_{2}=\beta(\cdot;-\frac{2\log(\epsilon_{1})}{\epsilon_{2}},f_{i}),\enspace f_{i}(x)=\frac{\log(x)-\boldsymbol{\mu}_{i}}{\left\lceil\frac{K}{100}\cdot N_{\text{eff}}\right\rceil\boldsymbol{\sigma}_{i}+\epsilon_{2}}$
where we note that $-\frac{2\log(\epsilon_{1})}{\epsilon_{2}}$ upper-bounds all the possible values that can be attained by $f_{i}$’s under our assumptions.
@@
$\kappa(\mathbf{X}^{\prime},\mathbf{p})=-\boldsymbol{\beta}_{2}(\mathbf{p})\quad T(\mathbf{X}^{\prime},\mathbf{p})=-\big{(}\text{sort}(\boldsymbol{\beta}_{2}(\mathbf{p}))_{\left\lceil\frac{K}{100}\cdot N_{\text{eff}}\right\rceil}\big{)}\quad g(\mathbf{X}^{\prime},\mathbf{p})=\frac{\boldsymbol{\beta}_{2}(\mathbf{p})}{\left\lceil\frac{K}{100}\cdot N_{\text{eff}}\right\rceil}$
@@
In most of our cases of interest, this follows from the fact that probabilities lie within $[0,1]$ and that padding is represented by $-1$. In other cases, e.g., the application of $\beta$, the sets might be different, but remain disjoint and compact.
@@
The finite disjoint union of compact subsets correspond to all possible sequence lengths ($N^{\prime}\in N_{\text{max}}$) and vocabulary sizes ($V^{\prime}\in V_{\text{max}}$). Below we provide the proof for Lemma A.3.
@@ B.1 Experimental Details
Our experiments were conducted using the PyTorch [36] framework (License: BSD), using a single NVIDIA L-40 GPU for all experiments regarding LOS-Net. We use a fixed batch size of 64 for all the tasks and datasets, and a fixed value of 8 heads (except for the Movies[35] dataset) in our light-weight transformer encoder for LOS-Net. Hyperparameter tuning was performed utilizing the Weight and Biases framework [5] – see Table 3.
@@ B.2 HyperParameters
In this section, we detail the hyperparameter search conducted for our experiments. We use the same hyperparameter grid for our main model, LOS-Net, and our proposed learning-based baselines, namely, ATP+R-MLP, ATP+R-Transf.. Additionally, we note that for a given dataset, we maintained the same grid search approach for all LLMs’ LOSs that we have trained on. The hyperparameter search configurations for all datasets are presented in Table 3. The grid search optimizes for the AUC calculated on the validation set.
@@ B.4 Our Baselines and Rank Encoding
ATP+R-Transf. This baseline is implemented as described in Equation 3, but without incorporating the TDS ($\mathbf{X}$), as follows:
@@
where $\mathcal{T}$ represents an encoder-only transformer architecture [50].
ATP+R-MLP. This baseline is similar to ATP+R-Transf. but replaces the transformer with an MLP. Formally:
@@ B.5.1 Datasets for Hallucination Detection
    Llama-3-8b-Instruct [46] (License: Llama-3777https://huggingface.co/meta-llama/Meta-Llama-3-8B/blob/main/LICENSE). Referred to as L-3-8b in the main text and accessed through the Hugging Face interface at https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct3.
@@ B.5.2 Datasets for Data Contamination Detection
B.6 Ablation Study
Existing methods often overlook a critical aspect of LOS. Specifically, they primarily rely on the ATP, $\mathbf{p}$, while neglecting the TDS, $\mathbf{X}$. In this subsection, we conduct an ablation study to evaluate the significance of the TDS in general, as well as its size, namely the hyperparameter $K$ introduced in Equation 1.

### v3

@@ 5.4 Run-Time Analysis
To empirically assess the efficiency of our approach, we ran a comprehensive set of training and inference timings, reported in full in Section˜B.10 and discussed in the following. The results clearly show LOS-Net features an extremely contained detection latency: $\sim 10^{-5}$s per inference fwd-pass. Training is also efficient, typically completing in under one hour on a single NVIDIA L-40 GPU, and often taking significantly shorter. To contextualize this computational efficiency w.r.t. methods relying on multiple prompting/generations [37, 25], we measured the detection latency of Semantic Entropy (SE) [13, 25], a pioneering method for HD. SE involves generating 10 additional responses and their semantic clustering, operated by checking mutual entailment with an auxiliary language model. On average, we measured $7.14\pm 2.97$ seconds per sample detection on Movies/L-3-8b and $7.55\pm 1.70$ seconds on Movies/Mis-7b. This is five orders of magnitude slower than LOS-Net, making the latter preferable for both accuracy and latency. Results on other LLM/dataset combinations are reported in Section˜B.10.
5.5 The value of $K$ and restricted TDS access
We conclude this section by presenting results on the impact of the parameter $K$, as defined in Equation˜1. In particular, we slide $K$ in $\{10,50,100,500,1000\}$ and discuss here results for the Mis-7B LLM on the HotpotQA dataset, see Table˜3. For each of the above values, we measure two quantities: the average probability mass captured, i.e., $\nicefrac{{\sum_{n=1}^{N}\left(\sum_{v=1}^{V}(\mathbf{X}^{\prime}_{n,v})\right)}}{{N}}$, and the corresponding performance of LOS-Net.
Table 3: Ablation study on $K$ in LOS-Net. Average Probability Mass (APM) and Test AUC for varying $K$ on Mis-7B - HotpotQA.
K APM (%) Test AUC
10 99.49 71.82 $\pm$ 0.15
50 99.80 71.87 $\pm$ 0.24
100 99.85 72.34 $\pm$ 0.20
500 99.99 72.67 $\pm$ 0.32
1000 99.99 72.92 $\pm$ 0.45
We observe that the former exceeds $0.99$ for all considered $K$’s, indicating that even values as small as $K=10$ are often sufficient to convey most of the information in the full TDS. In terms of performance, Test AUC tends to improve as $K$ values increase, though with diminishing returns beyond $K\geq 100$. As expected, the value $K=1000$ tends to deliver the best performance overall, this confirmed by results on the other LLM-dataset combinations reported in Section˜B.7. Given its extremely contained run-times (see above), this value appears to hit a sweet-spot optimizing performance and complexity. Most notably, however, even with $K=10$, LOS-Net outperforms all baselines and matches the white-box Activation Probe, highlighting the practical effectiveness of LOS-Net even in API-limited settings such as in GPT models – at the time of writing, exposing only the top $K=20$ output logits.
@@ 6 Conclusions
We proposed LOS-Net, an efficient method to detect data contamination and hallucinations in LLMs by leveraging their output signatures (LOS), defined as the union of Token Distribution Sequences (TDS) and Actual Token Probabilities (ATP). LOS-Net consists of a lightweight attention-based model operating on an effective encoding of the LOS. We proved it unifies and extends existing gray-box methods under a general framework, and experimentally showed it outperforms state-of-the-art gray-box methods across datasets and LLMs. It also exhibited promising generalization and transfer capabilities of LOS-Net, both across datasets and across LLMs. Our framework could be applied to other tasks, such as detecting LLM-generated content. Additional sources of information can also be incorporated, e.g., in the absence of latency constraints, it can be interesting to include “exact-token” flags as proposed by [37]. Last, the LOS can be extended to account for multiple prompting [25].
Acknowledgements
The authors are grateful to Beatrice Bevilacqua for insightful discussions. G.B. is supported by the Jacobs Qualcomm PhD Fellowship. F.F. conducted this work supported by an Aly Kaufman and an Andrew and Erna Finci Viterbi Post-Doctoral Fellowship. Y.G. is supported by the UKRI Engineering and Physical Sciences Research Council (EPSRC) CDT in Autonomous and Intelligent Machines and Systems (grant reference EP/S024050/1). H.M. is a Robert J. Shillman Fellow and is supported by the Israel Science Foundation through a personal grant (ISF 264/23) and an equipment grant (ISF 532/23). D.L. is funded by an NSF Graduate Fellowship. Research was also supported by the Israeli Ministry of Science, Israel-Singapore binational grant 207606. F.F. is extremely grateful to the members of the “Eva Project”, whose support he immensely appreciates.
@@ Appendix A Proofs
We define $\mathcal{D}\coloneqq\mathcal{X}\times\mathcal{M}.$ Recall that the target function we want to approximate is the gated scoring function $R$ as defined in Equation˜4, which can be written as follows:
@@
Define $f^{(1)}:\mathcal{D}\rightarrow\mathbb{R}^{N_{\text{max}}}$ to be the components of the sum in Equation˜5:
@@
As a result, the input to the transformer encoder in our architecture (see Equation˜2) becomes $\mathbf{X}^{\prime}||\mathbf{p}\in\mathbb{R}^{N_{\text{max}}\times(V_{\text{max}}+1)}$.
This simplifies our architecture in Equation˜2 to:
@@
$\left\lVert\mathcal{F}\right\rVert_{L^{1}}$ $=\int_{x\in\mathcal{D}}\left\lVert\mathcal{F}(x)\right\rVert_{1}\,dx=\int_{x\in\mathcal{D}}\sum_{i=1}^{n_{2}}{|\mathcal{F}(x)_{i}|}\,dx$
$=\sum_{i=1}^{n_{2}}\int_{x\in\mathcal{D}}|\mathcal{F}(x)_{i}|\,dx=\sum_{i=1}^{n_{2}}\left\lVert\mathcal{F}(x)_{i}\right\rVert_{L^{1}},$ (11)
@@
Next, observe that $f^{(1)}\in L^{1}(\mathcal{D})$. To see this, first note that $f^{(1)}$ is measurable. The indicator function is measurable because the indicator set is the preimage of the measurable function $\kappa(x)-T(x)$ on the closed set $0,\infty)$. Thus, $f^{(1)}$, being a product of measurable functions, is measurable. Next, we show that the $L^{1}$ norm is finite. This is true because $f^{(1)}$ is a product of the integrable function $g$ and the bounded function $\mathbf{1}$ on the compact domain $\mathcal{D}$.
@@
$\boldsymbol{\sigma}_{i}$ $=\sqrt{\mathbb{E}_{\mathbf{X}^{\prime}_{i}}[(\log(\mathbf{X}^{\prime}_{i})-\boldsymbol{\mu}_{i})^{2}]}$
$=\sqrt{\sum_{v=1}^{V}\mathbf{X}^{\prime}_{i,v}\cdot\big(\log(\mathbf{X}^{\prime}_{i,v})-\boldsymbol{\mu}_{i}\big)^{2}},$ (21)
Where $\mathbf{X}^{\prime}$ is given from Equation˜1.
@@
To prove Footnote˜4, it suffices to show the following. First (i), that the baselines can be implemented as in Equation˜4, given their sequence length and vocabulary size satisfy, $N\leq N_{\text{max}},\quad V\leq V_{\text{max}}$, where values in the inputs for indices larger than $N,V$ are ‘padded’ with e.g., 1-1. Second (ii), that their implementations are realized with $\kappa$, $T$, and $g$ which are all measurable, and with $g$ also integrable.
(i) Let us slightly modify the implementations provided in the Proof for Proposition˜1 to correctly account for padding values. Let us conveniently define:
$\alpha:\mathbb{R}\rightarrow\mathbb{R},$ $~~~\alpha(x)=1-\mathrm{ReLU}(-x)=\begin{cases}1&x\geq 0\\ 1+x&x<0\end{cases}$
@@
$\kappa(\mathbf{X}^{\prime},\mathbf{p})$ $=-\beta(\mathbf{p})$
$T(\mathbf{X}^{\prime},\mathbf{p})$ $=-\min(\beta(\mathbf{p}))$
$g(\mathbf{X}^{\prime},\mathbf{p})$ $=\mathbf{p}$
$M$ $=2,$
$f$ $\equiv\text{id.}$
@@
$\kappa(\mathbf{X}^{\prime},\mathbf{p})$ $=-\beta(\mathbf{p})$
$T(\mathbf{X}^{\prime},\mathbf{p})$ $=-\big(\text{sort}(\beta(\mathbf{p}))_{\left\lceil\frac{K}{100}\cdot N_{\text{eff}}\right\rceil}\big)$
$g(\mathbf{X}^{\prime},\mathbf{p})$ $=\frac{\log(\beta(\mathbf{p}))}{\left\lceil\frac{K}{100}\cdot N_{\text{eff}}\right\rceil}$
$M$ $=2,$
$f$ $\equiv\text{id.}$
@@
MinK%++. Before illustrating how this baseline is implemented, we note the following. In order for the normalization of log-probs to be well-defined, it is required that: (1) $\mathbf{\mu}$ is finite, (2) the denominator is greater than 0. As for (1), we note that null probability values ($X_{i,v}=0$) would be problematic, as they would cause the $\log$ function to output -$\infty$. We assume, in this case, that all probability values lie in $\epsilon_{1},1]$, with $\epsilon_{1}$ being a small value such that $0<\epsilon_{1}<1$. Regarding (2), we see that the problematic situation would occur in cases where the probability distribution is uniform. We assume to handle this case by adding a small positive constant $\epsilon_{2}>0$ in the denominator, so that the normalization would take the form: $\bar{\mathbf{p}}=\frac{\log(\mathbf{p})-\boldsymbol{\mu}}{\boldsymbol{\sigma}+\epsilon_{2}}$.
@@
$\beta_{1}$ $=\beta(\cdot;2,\text{id.})$
$\beta^{i}_{2}$ $=\beta(\cdot;-\frac{2\log(\epsilon_{1})}{\epsilon_{2}},f_{i}),$
$f_{i}(x)$ $=\frac{\log(x)-\boldsymbol{\mu}_{i}}{\left\lceil\frac{K}{100}\cdot N_{\text{eff}}\right\rceil\boldsymbol{\sigma}_{i}+\epsilon_{2}}$
where we note that $\frac{2\log(\epsilon_{1})}{\epsilon_{2}}$ upper-bounds all the possible values that can be attained by $f_{i}$’s under our assumptions.
@@
$\kappa(\mathbf{X}^{\prime},\mathbf{p})$ $=-\boldsymbol{\beta}_{2}(\mathbf{p})$
$T(\mathbf{X}^{\prime},\mathbf{p})$ $=-\big(\text{sort}(\boldsymbol{\beta}_{2}(\mathbf{p}))_{\left\lceil\frac{K}{100}\cdot N_{\text{eff}}\right\rceil}\big)$
$g(\mathbf{X}^{\prime},\mathbf{p})$ $=\frac{\boldsymbol{\beta}_{2}(\mathbf{p})}{\left\lceil\frac{K}{100}\cdot N_{\text{eff}}\right\rceil}$
@@
In most of our cases of interest, this follows from the fact that probabilities lie within $[0,1]$ and that padding is represented by 1-1. In other cases, e.g., the application of $\beta$, the sets might be different, but remain disjoint and compact.
@@
The finite disjoint union of compact subsets correspond to all possible sequence lengths ($N^{\prime}\in N_{\text{max}}$) and vocabulary sizes ($V^{\prime}\in V_{\text{max}}$). Below we provide the proof for Lemma˜2.
@@ B.1 Experimental Details
Our experiments were conducted using the PyTorch [38] framework (License: BSD), using a single NVIDIA L-40 GPU for all experiments regarding LOS-Net. We use a fixed batch size of 64 for all the tasks and datasets, and a fixed value of 8 heads (except for the Movies[37] dataset) in our light-weight transformer encoder for LOS-Net. Hyperparameter tuning was performed utilizing the Weight and Biases framework [6] – see Table˜4.
@@ B.2 HyperParameters
In this section, we detail the hyperparameter search conducted for our experiments. We use the same hyperparameter grid for our main model, LOS-Net, and our proposed learning-based baselines, namely, ATP+R-MLP, ATP+R-Transf.. Additionally, we note that for a given dataset, we maintained the same grid search approach for all LLMs’ LOSs that we have trained on. The hyperparameter search configurations for all datasets are presented in Table˜4. The grid search optimizes for the AUC calculated on the validation set.
@@ B.4 Our Baselines and Rank Encoding
Rank Encoding.  We construct the following learnable rank encoding  [^8] For the Wiki-MIA dataset, we used a lookup table for Rank Encoding, where the index corresponds to $r_{i}$ and the value is an embedding.,
Our baselines.  Below we present our additional learnable baselines. ATP+R-Transf is implemented as described in Equation˜2, but without incorporating the TDS ($\mathbf{X}$), as follows:
@@
where $\mathcal{T}$ represents an encoder-only transformer architecture [50]. ATP+R-MLP is similar to ATP+R-Transf. but replaces the transformer with an MLP. Formally:
@@ B.5.1 Datasets for Hallucination Detection
    Llama-3-8b-Instruct [47] (License: Llama-3999https://huggingface.co/meta-llama/Meta-Llama-3-8B/blob/main/LICENSE). Referred to as L-3-8b in the main text and accessed through the Hugging Face interface at https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct3.
@@ B.5.2 Datasets for Data Contamination Detection
B.6 Extended Transferability Experiments
Setup.  We fine-tune the LOS-Net given target LLM/dataset (depending on the task) for 10 epochs —significantly fewer than the number of epochs used in our standard training protocol. Notably, this fine-tuning process takes less than one minute in practice, on a single NVIDIA L-40 GPU.
To evaluate the effectiveness of fine-tuning, we benchmark the resulting model against two baselines. First, to assess knowledge transfer, we compare it with a LOS-Net model trained from scratch under the same 10-epoch setup. Second, we compare against the strongest known probability-based baselines with comparable detection latency—specifically, the strongest among Logits-mean/min/max and Probas-mean/min/max (see Table˜1). This comparison is essential: generalization scores above 0.5 AUC are only meaningful if they outperform non-learnable baselines that rely purely on probabilities or logits. We note that we exclude P(true) and Semantic Entropy baselines from this assessment, as they incur significantly higher latency (five order of magnitude higher than LOS-Net) and are thus not directly comparable to LOS-Net Ṫhese methods require additional generation or prompting to detect hallucinations. For a more detailed analysis, please refer to Section˜5.4.
Comprehensive results are shown in Figure˜4 (cross-LLM generalization) and Figure˜5 (cross-dataset generalization).
In the heatmaps, a superscript ‘’ indicates that the fine-tuned LOS-Net outperforms its scratch-trained counterpart in the same setting—evidence of successful transfer. Bold entries denote cases where it surpasses the best non-learnable probability-based method.
B.7 Ablation Study
Existing methods often overlook a critical aspect of LOS. Specifically, they primarily rely on the ATP, $\mathbf{p}$, while neglecting the TDS, $\mathbf{X}$. In this subsection, we conduct an ablation study to evaluate the significance of the TDS in general, as well as its size, namely the hyperparameter $K$ introduced in Equation˜1.